# thanks: current affiliation: IBM Quantum

### v1

@@ I.2 Logical dynamical decoupling
grouping together terms with similar action on the code space $\mathcal{C}$. Altogether, this expression has $4^{n}$ linearly independent terms corresponding to the elements of $\tilde{\mathcal{P}}_{n}$. Specifically, $H_{SB}^{\mathcal{S}}$ collects all the distinct $2^{n-k}$ Pauli operators $P_{i}\in\mathcal{S}$ (including $P_{0}=I$); such terms act as the identity operator on the code space and, therefore, are harmless. $H_{SB}^{\mathcal{L}}$ collects all $2^{n-k}(4^{k}-1)$ Pauli operators in $(\mathcal{S}\mathcal{L}\setminus\mathcal{S})/\{\pm 1,\pm i\}$; these terms leave the code space (and each of the other syndrome spaces) invariant but cause harmful logical errors, since by construction, the code cannot detect such errors. Finally, $H_{SB}^{\mathcal{E}}$ collects all remaining $2^{n+k}(2^{n-k}-1)$ terms in $\tilde{\mathcal{P}}_{n}\setminus(\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\})$, which correspond to potentially detectable errors: given the value of the syndrome one can either detect all of them or correct some of them.
A key observation is that the decoupling group is arbitrary and may, in particular, be formed from the stabilizer group $\mathcal{S}$ and/or the group of canonical logical operators $\mathcal{L}$ of a stabilizer code $\mathcal{C}$, or even a subset of either one [87, 72, 88, 89, 58, 90]. One approach for selectively combining DD with QEC, which leverages this flexibility, is to use DD only to suppress the errors between or within the $2^{n-k}$ syndrome spaces, including the code space $\mathcal{C}$ (i.e., errors in $(\mathcal{S}\mathcal{L}\setminus\mathcal{S})/\{\pm 1,\pm i\}$). Errors causing transitions between the syndrome spaces (i.e., elements of $\tilde{\mathcal{P}}_{n}\setminus(\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\})$) are the ones the code aims to detect or correct. The former are decoupled by $\mathcal{G}=\mathcal{L}$ while the latter are decoupled by $\mathcal{G}=\mathcal{S}$. For any choice of $\mathcal{G}$, the decoupling condition is that there exists at least one element of $\mathcal{G}$ that anticommutes with a given error. Based on these properties, it was shown in Ref. [58] that combining the full set of $n+k$ stabilizer and canonical logical operator generators to generate the DD group $\mathcal{G}=\mathcal{S}\times\mathcal{L}$, optimally achieves the goal of decoupling the errors between or within the $2^{n-k}$ syndrome spaces, in the sense of minimizing the length of the DD pulse sequence. This is known as ‘stabilizer-logical DD’ (SLDD) and incurs a single sequence cost of $T=2^{n+k}\tau$, instead of $4^{n}\tau$ in the case of full Pauli group decoupling. The savings are substantial when $k\ll n$. Adding QEC syndrome measurements and recovery operations based on syndrome measurements then detects the remaining errors between syndrome spaces that DD did not suppress [i.e., the $O(T^{q+1})$ terms in the case of a $q$’th-order DD sequence] or successfully corrects some of them, resulting in a hybrid QEC-DD strategy.
Figure 1: (a) Schematic of the experimental design: we encode a logical Bell state $\chi$ using the $[[4,2,2]]$ code and let it undergo either free evolution or LDD for a duration $\tau$. We then run the unencoding circuit for $\chi^{\prime}$ and measure the resulting bitstring. The encoding and unencoding circuits together take up to $5\mu$s. (b) The logical error set that relates the four Bell states as a result of a logical operator acting on one ($i=1$ or $2$) of the two logical qubits. (c) Interpretation of different bitstrings for an encoded state ($U_{\rm{enc},\chi}$), depending on the unencoding ($U_{\rm{enc},\chi^{\prime}}^{\dagger}$) circuit used. The color code indicates that the encoded state has undergone either no error (black) or one of the logical errors (red, blue, green) indicated in (b). The DD sequences used in our experiments are (d) Robust Logical XX (RLXX), (e) Robust Logical XY4 (RLXY4), and (f) physically Staggered XY4 (SXY4). Each panel shows both the physical implementation (top) and the logical interpretation of each sequence. A tilde in (d) and (e) denotes a negative $\pi$ rotation, i.e., $X=R_{X}(\pi)=\exp(-i\pi\sigma^{x})$ $\tilde{X}=R_{X}(-\pi)=\exp(i\pi\sigma^{x})$, and likewise for $Y$.
II The QEC-LDD Theorem
@@
The main weakness of SLDD is that it still incurs a cost that is exponential in the number of physical qubits $n$. Here, we present a different hybrid approach, shifting more of the burden for handling correctable errors to the QEC code and using only the logical operators as DD pulses $g_{j}$. We call this approach ‘logical DD’ (LDD). Compared to SLDD, this approach has the desired effect of reducing the single sequence time-cost to $T=4^{k}\tau$ for first-order decoupling, independent of $n$. Table 1 summarizes the relative resource cost of full Pauli group DD, SLDD, and LDD.
@@
Decoupling group $\mathcal{G}$ of pulses suppresses all errors afflicting code space / requires QEC
Pauli group $\tilde{\mathcal{P}}_{n}$ $2^{2n}$ Yes / No
SLDD $2^{n+k}$ Yes / No
LDD $2^{2k}$ No / Yes
@@
Table 1: The comparative resource cost for various DD methods in terms of the number of pulses required to suppress the most general system-bath Hamiltonian $H_{SB}$ [Eq. 5] acting on $n$ qubits, and whether active error correction is required in addition. Full Pauli group DD and SLDD suppress all error mechanisms which can decohere code states, while LDD does not, as it leaves dealing with these error mechanisms for the QEC cycle. Consequently, LDD does not scale with $n$ but with $k$, whereas Pauli group DD and SLDD both scale with $n$.
II.1 Theorem statement and proof
The LDD sequence is purely responsible for decoupling the logical errors acting within each syndrome space, but unlike SLDD, it no longer decouples some of the physical errors between the syndrome spaces. Correcting these errors is accomplished by syndrome measurements in conjunction with error recovery.
For each code $\mathcal{C}$ there is a choice of which errors to correct: for each of the $12^{n-k}-$ 1 nontrivial syndromes, one may choose a recovery operation; this choice determines which errors are successfully corrected and which result in a logical error. Once such a choice is made, we call the errors that are successfully corrected “correctable”.
We summarize this in the following theorem:
Theorem 1 (QEC-LDD).
Consider the most general system-bath Hamiltonian acting on $n$ qubits [Eq. 5]. Let $\mathcal{C}$ be an $[[n,k,d]]$ stabilizer code (with any $d\geq 1$). Let $\mathcal{L}=\langle i,\overline{X}_{j},\overline{Z}_{j}\rangle_{j=1}^{k}$ denote its group of canonical logical operators. Choose the decoupling group $\tilde{\mathcal{G}}=\mathcal{L}/\{\pm 1,\pm i\}$, henceforth the LDD group. Then $\tilde{\mathcal{G}}$ decouples all logical errors, that is, every term in $H_{SB}^{\mathcal{L}}$, as well as some errors in $H_{SB}^{\mathcal{E}}$. There is a choice of recovery operation such that all remaining non-decoupled errors in $H_{SB}^{\mathcal{E}}$ are correctable by $\mathcal{C}$. Conversely, for any recovery operation chosen in this way, only those errors that are not decoupled by $\tilde{\mathcal{G}}$ are correctable.
Proof.
We first establish when an error $E\in\mathcal{P}_{n}$ is decoupled. Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ be the Hilbert space of $n$ qubits. Then $\mathcal{C}\subset\mathcal{H}$ and all elements of $\mathcal{P}_{n}$ (and, hence, of $\mathcal{L}\subset\mathcal{P}_{n}$ and $\mathcal{S}\subset\mathcal{P}_{n}$) are unitary operators on $\mathcal{H}$. We let $\mathcal{G}=\mathcal{L}=\{\pm g_{j},\pm ig_{j}\}_{j=0}^{4^{k}-1}$. Consider a term in $H_{SB}$ corresponding to some $E\in\mathcal{P}_{n}$. It is decoupled if and only if there exists $g\in\mathcal{G}$ anticommuting with $E$. Indeed, then $g^{\dagger}Eg=-E$ and
$\langle E\rangle_{\mathcal{G}}=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}h^{\dagger}Eh=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}(gh)^{\dagger}E{gh}=-\langle E\rangle_{\mathcal{G}},$ (6)
so $\langle E\rangle_{\mathcal{G}}=0$. If there is no such $g$ then all elements of $\mathcal{G}$ commute with $E$ (because every two elements of $\mathcal{P}_{n}$ either commute or anticommute). Then all $g^{\dagger}Eg=E$ for all $g\in\mathcal{G}$, hence $\langle E\rangle_{\mathcal{G}}=E$ (i.e., the term is undecoupled). It follows that $\tilde{\mathcal{G}}$ decouples all logical errors, since every such error anticommutes with an element of $\mathcal{G}$. The errors in $H_{SB}^{\mathcal{E}}$ that anticommute with an element of $\mathcal{G}$ are also decoupled. All terms in $H_{SB}^{\mathcal{S}}$ are undecoupled since they all commute with $\mathcal{G}$.
The key observation in the proof is that the statement of the theorem is invariant under Clifford maps, which we now show.
Now consider any Clifford unitary $V\colon\mathcal{H}\to\mathcal{H}$. Let $\mathcal{C}^{\prime}=V\mathcal{C}$, $d^{\prime}\geq 1$ be the distance of the code $\mathcal{C}^{\prime}$, $\mathcal{G}^{\prime}=\mathcal{L}^{\prime}=V\mathcal{L}V^{\dagger}$, $\tilde{\mathcal{G}}^{\prime}=\mathcal{G}^{\prime}/\{\pm 1,\pm i\}$, $H_{SB}^{\prime}=VH_{SB}V^{\dagger}$, etc. For a recovery map $R_{s}$ (for every value $s$ of the syndrome) for $\mathcal{C}$ we can construct the corresponding recovery map $R_{s}^{\prime}=VR_{s}V^{\dagger}$ for $\mathcal{C}^{\prime}$. If the statement of the theorem is true for the original code $\mathcal{C}$, then it remains true for $\mathcal{C}^{\prime}$ because it only depended on the commutation relations between elements of $\mathcal{P}_{n}$ and those are preserved under the group homomorphism $P\mapsto VPV^{\dagger}$. By applying this observation to $V^{\dagger}$ instead of $V$ we observe that the converse also holds.
In 1 in Appendix C, extending earlier methods for stabilizer code encoding [91, 92, 93], we show that for any code $\mathcal{C}$ and any choice of the canonical logical operators $\mathcal{L}$ there is a Clifford unitary $V$ (which we may call “unencoding circuit”) such that the corresponding code $\mathcal{C}^{\prime}$ consists of states of the form |$\ket{\psi}\otimes\ket{0}^{n-k}$, where |$\ket{\psi}\in(\mathbb{C}^{2})^{\otimes k}$, and $\mathcal{L}^{\prime}=V\mathcal{L}V^{\dagger}\simeq\mathcal{P}_{k}$ consists of Paulis $P\in\mathcal{P}_{n}$ acting trivially on the last $n-k$ qubits. We call such code $\mathcal{C}^{\prime}$ a trivial code: its logical qubits can be chosen to coincide with the first $k$ physical qubits and stabilizer generators can be chosen to be $Z_{k+1},\dots,Z_{n}$ enforcing that the last $n-k$ qubits are set to 00, and $d^{\prime}=1$.
It follows that it is sufficient to prove the statement of the theorem for this trivial code. Hence, from now on we assume that $\mathcal{C}$ is the trivial code described above. $E\in\mathcal{P}_{n}$ is undecoupled if and only if it commutes with all Paulis acting on the first $k$ qubits. This is true if and only if $E$ acts trivially on the first $k$ qubits. (Equivalently, $E\in\mathcal{P}_{n}$ is decoupled if and only if acts non-trivially on one or more of the first $k$ qubits.) Measuring the stabilizers $Z_{k+1},\dots,Z_{n}$ is equivalent to measuring the last $n-k$ qubits. Any such measurement outcome $s=(s_{1},\dots,s_{n-k})$ corresponds to errors $E=SE_{s}$, where $E_{s}=\prod_{j=1}^{n-k}X_{k+j}^{s_{j}}$ and $S$ is an arbitrary stabilizer. A (unitary) recovery operation $R_{s}$ corrects such an error $E$ if and only if $R_{s}=S^{\prime}E_{s}$. In particular, if we apply $E_{s}$ corresponding to the measured syndrome $s$, we correct all non-decoupled errors, proving the first statement of the theorem. On the other hand, all such recovery maps act trivially on the first $k$ qubits, so if $E$ acts non-trivially on the first $k$ qubits (i.e. if $E$ is decoupled), it remains uncorrected by any such procedure. ∎
We emphasize a few key points:
  • The claim holds even for $d=2$ (or $d=1$), i.e., an error detecting code becomes an error correcting code subject to LDD.
  • In fact, the notion of correctable errors used here ignores the code distance altogether. For any distance, the QEC-LDD protocol has at least $3k$ non-correctable single-qubit errors (these are decoupled by LDD). [^1] Consider the group $\mathcal{L}$ and divide the physical qubits into three classes according to the action of the elements of $\mathcal{L}$ on a particular qubit. 1. Every element of $\mathcal{L}$ acts as $I$ on that qubit. 2. Every element of $\mathcal{L}$ acts as $I$ or one other Pauli ($X$, $Y$, or $Z$) on that qubit. 3. For every Pauli in $\{I,X,Y,Z\}$ there is $P\in\mathcal{L}$ which acts like that Pauli on that qubit. Let $l_{1},l_{2},l_{3}$ be the number of physical qubits in each class. Restrict every element of $\mathcal{L}$ to the qubits from class 3. Such a restriction does not change the commutation relations. Let the resulting group be $\mathcal{L}^{\prime}$. We have $\mathcal{P}_{k}\simeq\mathcal{L}\simeq\mathcal{L}^{\prime}$, which in turn is embedded in $\mathcal{P}_{l_{3}}$. Therefore, $l_{3}\geq k$. The number of uncorrectable single-qubit errors is $2l_{2}+3l_{3}\geq l_{3}\geq 3k$. E.g., for the $[[4,2,2]]$ code $l_{2}=0,l_{3}=3,k=2$, $2l_{2}+3l_{3}=9>6=3k$. For example, in the standard distance-$d$ surface code there is a weight $1$ error, which can be “corrected” to become a weight $12d-$ 1 error. That does not contradict the theorem, because that error is suppressed by LDD.
  • For each syndrome, the corresponding recovery operation is unique up to multiplication by stabilizers.
  • All decoupled errors are not correctable; it is the role of LDD to suppress such errors.
II.2 Illustration of the QEC-LDD Theorem: the $[[4,2,2]]$ code

### v2

@@ II.2 Dynamical decoupling and quantum error correction
Eq. 6 groups terms with similar action on the code space $\mathcal{C}$. In total, $H_{SB}+I_{S}\otimes H_{B}$ has $4^{n}$ linearly independent terms corresponding to the elements of $\tilde{\mathcal{P}}_{n}$. Specifically, $H_{SB}^{\mathcal{S}_{\star}}+I_{S}\otimes H_{B}$ collects all $2^{n-k}$ distinct terms in $\{I_{S}\}\dot{\cup}\mathcal{S}_{\star}$ that act as the identity operator on the code space and, therefore, are harmless. $H_{SB}^{\mathcal{L}_{\star}}$ collects all $2^{n-k}(4^{k}-1)$ Pauli operators in $\mathcal{L}_{\star}$; these terms are non-trivial logical operators that leave the code space (and each of the other syndrome spaces) invariant. Finally, $H_{SB}^{\mathcal{D}_{\star}}$ collects all remaining $4^{n}-2^{n-k}-2^{n-k}(4^{k}-1)=4^{n}-2^{n+k}$ terms in $\mathcal{D}_{\star}$, which correspond to detectable and potentially correctable) errors.
We can now establish when an error $E\in\mathcal{P}_{n}$ is decoupled.
Lemma 1.
Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ be the Hilbert space of $n$ qubits. Then all elements of $\mathcal{P}_{n}$ are unitary operators on $\mathcal{H}$. Consider a term in $H_{SB}$ corresponding to some $E\in\mathcal{P}_{n}$. It is decoupled by $\tilde{\mathcal{G}}$ in the sense of 1 if and only if there exists $g\in\mathcal{G}$ anticommuting with $E$.
Proof.
If there exists $g\in\mathcal{G}$ that anticommutes with $E$ then $g^{\dagger}Eg=-E$ and
$\langle E\rangle_{\mathcal{G}}=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}h^{\dagger}Eh=-\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}(gh)^{\dagger}E{gh}=-\langle E\rangle_{\mathcal{G}},$ (7)
so the group average $\langle E\rangle_{\mathcal{G}}=0$. If there is no such $g$, then all elements of $\mathcal{G}$ commute with $E$ (because every two elements of $\mathcal{P}_{n}$ either commute or anticommute). Then $g^{\dagger}Eg=E$ for all $g\in\mathcal{G}$, hence $\langle E\rangle_{\mathcal{G}}=E$ (i.e., the term is undecoupled). ∎
II.3 Encoded dynamical decoupling
A key observation is that the decoupling group is arbitrary [70, 69] and may, in particular, be formed from the stabilizer group $\mathcal{S}$ and/or the group of logical operators $\mathcal{L}$ of a stabilizer code $\mathcal{C}$ [58]. We call these different choices “encoded dynamical decoupling,” as they all involve the use of encoded operations (relative to a QEC code) as DD pulses, instead of physical DD pulses. We next provide several examples of encoded DD groups.
When using a QEC code, the approach of using the full Pauli group $\mathcal{P}_{n}$ is excessive, since there is no need to decouple the terms in $H_{SB}^{\mathcal{S}_{\star}}$; this observation motivated Ref. [58] to introduce SLDD, which achieves a reduction from a DD group order of $4^{n}$ to $2^{n+k}$. The savings are substantial when $k\ll n$.
Definition 2 (SLDD).
For any stabilizer group $\mathcal{S}$ and associated canonical group of logical operators $\mathcal{L}$, the stabilizer-logical dynamical decoupling (SLDD) group is
$\tilde{\mathcal{G}}_{\text{SLDD}}=\pi(\mathcal{S}\mathcal{L}).$ (8)
Note that $\mathcal{S}\mathcal{L}=\mathcal{N}(\mathcal{S})$ and thus independent of the choice of $\mathcal{L}$.
@@ II.3 Encoded dynamical decoupling
The main weakness of SLDD is that it still incurs a cost that is exponential in the number of physical qubits $n$. This observation motivates us to consider a different hybrid approach that shifts more of the burden for handling correctable errors to the QEC code and uses only the logical operators of the code as DD pulses:
Definition 3 (LDD).
Pick a canonical group of logical operators $\mathcal{L}$. The logical dynamical decoupling (LDD) group is
$\tilde{\mathcal{G}}_{\text{LDD}}\equiv\tilde{\mathcal{L}}=\pi(\mathcal{L}).$ (9)
Compared to SLDD, this approach has the desired effect of reducing the DD group order to $T\leq 4^{k}$, independent of $n$.
For the sake of completeness, we also mention the case of DD using only the stabilizers:
Definition 4 (SDD).
For any stabilizer group $\mathcal{S}$, the SDD group is
$\tilde{\mathcal{G}}_{\text{SDD}}\equiv\tilde{\mathcal{S}}=\pi(\mathcal{S}).$ (10)
SDD can be seen as an alternative to QEC since it decouples all detectable errors. However, it does not decouple any logical errors and is therefore not a particularly useful method when used in conjunction with error correction. It does play an important role in scenarios where QEC is not viable, such as in adiabatic quantum computing [79, 80].
We can now classify the different decoupling groups by the errors they decouple.
Theorem 1 (Decoupling by group type).
  1. 1.
    SDD decouples all detectable errors: $\langle{H_{SB}^{\mathcal{D}_{\star}}}\rangle_{\tilde{\mathcal{G}}_{\text{SDD}}}=0$.
  2. 2.
    LDD decouples all logical errors: $\langle{H_{SB}^{\mathcal{L}_{\star}}}\rangle_{\tilde{\mathcal{G}}_{\text{LDD}}}=0$.
  3. 3.
    SLDD decouples all errors: $\langle{H_{SB}^{\mathcal{L}_{\star}}+H_{SB}^{\mathcal{D}_{\star}}}\rangle_{\tilde{\mathcal{G}}_{\text{SLDD}}}=0$.
Proof.
It follows from 1 that it suffices to show that for each term in $H_{SB}^{\mathcal{S}_{\star}}$, $H_{SB}^{\mathcal{L}_{\star}}$, or $H_{SB}^{\mathcal{D}_{\star}}$, there exists $g\in\mathcal{G}$ it anticommutes with, where $\mathcal{G}$ is the corresponding decoupling group. Indeed:
  1. 1.
    SDD: Each detectable error anticommmutes with at least one element of $\mathcal{S}$.
  2. 2.
    LDD: Each nontrivial logical error anticommmutes with at least one element of $\mathcal{L}$.
  3. 3.
    Since $I\in\mathcal{L}$, the stabilizer $\mathcal{S}\subset\mathcal{G}_{\text{SLDD}}=\bigl{\{}\mathcal{S}L\mid L\in\mathcal{L}\bigr{\}}$. Thus, the group average of $H_{SB}^{\mathcal{L}_{\star}}+H_{SB}^{\mathcal{D}_{\star}}$ with respect to $\mathcal{G}_{\text{SLDD}}$ includes both $\mathcal{S}$, which decouples all detectable errors, and $\mathcal{L}$, which decouples all logical errors.
Table 1 summarizes the relative resource cost of full Pauli group DD, SDD, LDD, and SLDD.
@@
Decoupling group $\mathcal{G}$ of pulses error type suppressed requires QEC
$\tilde{\mathcal{P}}_{n}$ $2^{2n}$ all No
SDD $2^{n-k}$ detectable No
LDD $2^{2k}$ logical Yes
SLDD $2^{n+k}$ detectable & logical No
@@
Table 1: The comparative resource cost for various DD methods in terms of the number of pulses required to suppress the most general system-bath Hamiltonian $H_{SB}$ [Eq. 6] acting on $n$ qubits, and whether active error correction is required in addition. Full Pauli group DD and SLDD suppress all error mechanisms that can decohere code states, while LDD does not, as it leaves dealing with these error mechanisms for the QEC cycle. Consequently, LDD does not scale with $n$ but with $k$, whereas Pauli group DD and SLDD both scale with $n$.
It is common to use only subgroups of the complete group of logical operators. In the context of standard (unencoded) DD, this is the case, e.g., when using single-axis sequences such as XX instead of the universal XY4 sequence. In the present context, we would like to consider subgroups of $\mathcal{N}(\mathcal{S})$ other than $\mathcal{S}$ and $\mathcal{L}$, which will similarly suppress selected subsets of errors. We refer to this as normalizer DD, since the corresponding DD group elements are generic normalizer elements, as opposed to canonical logical operators.
Definition 5 (NDD).
Normalizer DD (NDD) is any pulse sequence formed from using a subgroup of $\tilde{\mathcal{G}}_{\text{SLDD}}$ other than $\tilde{\mathcal{G}}_{\text{SDD}}$ or $\tilde{\mathcal{G}}_{\text{LDD}}$.
NDD arises in our experiments, as described in Section III. We also discuss an example in the next subsection.
It is important to emphasize that additional errors may and will generally be decoupled by each of the decoupling groups, beyond those mentioned in 1. For example, a detectable weight-$1$ Pauli operator that anticommutes with an element of $\mathcal{G}_{\text{LDD}}$ will be decoupled despite not being a logical error for any code with $d\geq 2$.
Another important point is that since we are employing a Hamiltonian noise model, any terms that arise in second order perturbation theory, i.e., the $O(T^{2})$ terms we are neglecting in Eq. 2, will mix $H_{SB}^{\mathcal{S}_{\star}}$, $H_{SB}^{\mathcal{L}_{\star}}$, and $H_{SB}^{\mathcal{D}_{\star}}$ through commutators. This means that the clean separation we have assumed between error types disappears at this higher order of perturbation theory. However, higher-order decoupling sequences are known that achieve suppression up to arbitrary order $q$ [i.e., leaving an $O(T^{q+1})$ unsuppressed error term in $U(T)$[72, 81, 82, 83, 84]. For simplicity, in this work we restrict our attention to first-order sequences.
II.4 Illustration of LDD and NDD: the $[[4,2,2]]$ code
To illustrate the concepts discussed above, we analyze the example of the $[[4,2,2]]$ code, as it plays a key role in our experiments.